The elementary chemical reaction O + CIO >Cl + O2 is made pseudo-first order in oxygen atoms by using a large excess of CIO radicals. The rate constant for the reaction is 3.8 cm^3/molecule s. If the initial concentration of CIO is 1.0 x 10^11 molecules/cm', how long will it take for the oxygen atoms to decrease to 10.% of their initial concentration?
1
Expert's answer
2019-12-27T01:04:54-0500

t=ln(C0/C)/K=ln(1×1011/1×1010)/3.8=0.606 (s)

Answer: 0.606 s.